Hello hope you are having a good day
I am doing simplifying radicles in school and am have trouble with the following question
( radical sign ) 384 x to the power of 4 y cube

Answer by stanbon(75887)

( radical sign ) 384 x to the power of 4 y cube
--------
= sqrt[384x^4y^3]
-----
Note:: 384 = 2*192 = 2*2*96 = 2*2*2*48 = 2*2*2*2^4*3 = 2^7*3
----
= sqrt[2^7*3*x^4*y^3]
----
= 2^3*x^2*y*sqrt[2*3*y]
----
= 8x^2*y*sqrt[6y]
-------
Cheers,
Stan H.
-----------